# Exploring Community-Powered Conversational Agent for Health Knowledge Acquisition: A Case Study in Colorectal Cancer

### v1

@@ 3.2. Design of Potential Features
Our design is informed by Kuhlthau’s Information Search Process (ISP) model (Case and Given, 2016; Kuhlthau, 1991), which has been widely applied in educational contexts to understand how individuals seek and process information when faced with knowledge gaps. It outlines six stages that users typically go through in learning tasks: initiation, selection, exploration, formulation, collection, and presentation (Kuhlthau, 1991). Based on this model and prior literature in the education domain, we designed six potential features to support users in learning health knowledge using data from online communities. For example, in the “Selection” stage of the ISP model, users are expected to identify and choose a general topic for investigation. However, grounded in the Xie and Cool (2009)’s research, individuals often experience difficulty initiating the search process due to uncertainty or lack of clarity about the topic. To address this, and inspired by the chunking mechanism in learning (Gobet et al., 2001), we designed a feature—Switch example questions about other topics—which categorizes community data into distinct topics, and we expected it could enable users to explore different directions and more easily select a focus for further learning. Similarly, in alignment with the “Exploration” stage, users are expected to investigate the content of a general topic to deepen their personal understanding. However, they may encounter difficulties in articulating their information needs precisely (Kuhlthau, 1991; Xie and Cool, 2009). Inspired by prior tools that leverage LLMs to provide recommended questions, we designed the feature “Provide suggested follow-up questions” to support users in formulating relevant queries more easily during this process. Table 2 shows the six potential features, and the screenshots of these features are attached in Appendix A.
@@ 3.4. Findings
Provide suggested follow-up questions (Xie and Cool, 2009; Li et al., 2024a; Peng et al., 2024; Wu et al., 2024) 6.01 1.00
Switch example questions about other topics (Xie and Cool, 2009; Gobet et al., 2001) 5.96 1.01
Offer input suggestions when the user enters the keyword (Xie and Cool, 2009; Wu et al., 2024) 5.86 1.18
Provide real-world cases for reference (Xie and Cool, 2009; Lieb and Goel, 2024) 5.75 1.30
Assume a scenario to ask for an explanation (Sabnis and Nagashima, 2024; Lieb and Goel, 2024; Pan et al., 2025) 4.84 1.53
Test learning feedback with single-choice question (Li et al., 2024b; Lieb and Goel, 2024; Peng et al., 2024; Ruan et al., 2019) 4.64 1.87
We use the reflexive thematic analysis method (Braun and Clarke, 2012) to analyze the open-ended responses from the survey study as well as the transcribed recordings of each participant’s semi-structured interviews and co-design sessions.
@@ 3.5. Design Goals
DG1: The CA should ground its response to the user query on professional documents (Fan et al., 2024; Wang et al., 2024; Jin et al., 2016). Authority is the first priority for knowledge acquisition in the health domain. Many online information, while easily accessible, could be unauthoritative (C1). We should prepare professional documents that are validated by experts (e.g., certified doctors) as the base for the tool to provide responses to the user query.
DG2: The CA should assist users in exploring multiple knowledge aspects of the disease (Czeresnia, 1999; Wicks and Frost, 2008), such as the cause, symbols, diagnosis, and treatment. This design goal highlights the breadth of knowledge coverage, aiming to help users develop a holistic understanding of the disease. The rich information online can cover every knowledge aspect, but such information is usually unstructured and non-personalized (C1). Categorizing frequently asked questions into topics could support users in exploring the knowledge aspects of their interests, and this feature was also anticipated by our participants ($M=5.96$).
DG3: The CA should provide real-world cases of others’ experiences related to the user’s (Mamykina et al., 2015; Young and Miller, 2019; Willis et al., 2023). Participants went to online communities like RED to gain health information from others’ experiences. They can gain emotional support at the same time, which could reduce their anxiety (C2). The perceived usefulness of this feature was also highly rated by our participants ($M=5.75$).
DG4: The CA should assist users in delving into the current knowledge aspect of the disease (Czeresnia, 1999; Wicks and Frost, 2008). That is, it should support users in deepening their understanding of a specific topic they are currently exploring. This goal emphasizes the depth of knowledge acquisition and sustained engagement within a focused knowledge thread. Participants tended to give up seeking information online if they lacked support when confusion arose (C2). Suggesting follow-up questions ($M=6.01$) or input suggestions ($M=5.86$) could help them proceed to the learning process with CanAnswer, which were also favored by the participants in the study (Table 2).
@@ 5.1.1. Leveraging Doctor-Led Community Data
To satisfy DG1 and provide reliable, expert-verified responses, we first incorporate professional documents. Insights from our formative study suggest that tools designed to support health-related learning must ground their responses in trustworthy knowledge sources (Fan et al., 2024; Wang et al., 2024; Jin et al., 2016). One potential approach is to curate a base dataset of professional documents (e.g., clinical guidelines, academic literature) with the assistance of domain experts. While such manually curated datasets can provide high-level guidance—such as helping determine appropriate topic categories for topic switching (DG4)—they are often insufficient in scope and depth for directly supporting large language models (LLMs) in generating trustworthy and detailed answers.
@@ 5.2.1. The Base Dataset of Professional Documents
To start with, we closely collaborated with a professional medical team to compile a database of clinical question-answering (QA) pairs related to colorectal cancer. First, we distributed a questionnaire to 68 staff in a local governmental project for screening colorectal cancer to collect the related questions commonly asked by people. We identified 16 frequently asked questions from the collected data. Second, we analyzed 100 publicly available doctor-patient conversation records from the Dingxiang Doctor, using the “colorectal cancer” and “colonoscopy” as the search keywords. This analysis yielded an additional 14 common questions. Third, we prompt GPT-4 to simulate queries by inputting the prompt, “If you were a first-degree relative of a colorectal cancer patient, what would you want to know about colorectal cancer and colonoscopy? ” This simulation identified one additional question that was not covered in the previous two sources. After further discussion within the medical team, our final base professional documents consist of 24 questions. The medical team compiles answers to these questions by referring to fifteen professional guidebooks  [^10] Full list of the guidebooks are provided in the supplementary materials. (e.g., “Asia Pacific Consensus recommendations for colorectal cancer screening” (Sung et al., 2008)). To ensure the reliability of the compiled QA pairs, we consulted a panel of experts, including one physician specializing in cancer screening, two chief nurses with extensive experience in colorectal cancer care and colonoscopy screening, and five researchers specializing in cancer screening and disease prevention and control. The resulting 24 clinical QA pairs serve as the base dataset of professional documents in CanAnswer.
@@ 5.2.5. Topic switch
We utilize the 24 QA pairs in the Base Dataset ( Section 5.2.1) to group the topics related to colorectal cancers. The Enriched Dataset ( Section 5.2.2) was excluded from this step because it contains content, while relevant, that is not specifically about colorectal cancer. Given that the topics here aim to provide a broad, guideline-level overview specifically for learning about colorectal cancer, we opted to use only the Base Dataset. The initial categorization criteria were generated using LLOOM (Lam et al., 2024), powered by LLMs, which creates 16 topics along with the criteria for classification. These topics were then reviewed and named by an expert. The finalized list of topics includes: Medical Definitions, Medication Use, Medical Decision-Making, Colon Cancer Treatment, Symptoms and Signs, Misconceptions About Colonoscopy, Family History Risks, Cancer Prevention, Dietary Focus, Risk Factor Identification, Pain Management, Follow-Up Care, Infection Prevention, Dietary Preparation, Colonoscopy Information, and Colon Cancer Screening Guidelines. When users interact with the topic switch dropdown menu, our system classifies their most recent query using GPT-4o, guided by refined standards generated by LLOOM. This mechanism accurately matches the user’s latest question to the most relevant topic in 16 topics. Once the user selects a new topic, CanAnswer prompts GPT-4o to generate follow-up questions by utilizing the chosen topic and the user’s chat history.
@@ 5.2.6. Processing Peer-led Online Community Data for CanAnswer
Our system recommends community posts from RED to provide users with relevant real-world cases related to colorectal cancer. We leveraged an open-source GitHub project  [^17] https://github.com/NanmiCoder/MediaCrawler to collect an initial dataset of 1,961 posts using keywords such as “intestinal polyps”, “hematochezia”, “intestinal cancer”, “chemotherapy regimen for colorectal cancer”, “immunotherapy for colorectal cancer”, “surgical resection for colorectal cancer”, “intestinal obstruction”, “colonoscopy”, “enteritis”, “flatulence”, and “diarrhea”. Then, we use tags associated with each post to classify these posts, as these tags were added by the post creators and could reflect the intentions and targeted audience of the posts. Specifically, these tags were input into GPT-4o, which generated a preliminary set of categories. The categories were iteratively refined through six rounds of feedback and manual verifications by two researchers, prioritizing the specificity of conditions and disease types to improve relevance. An expert (doctor with 30 years of clinical experience) then reviewed the final classification to validate its appropriateness. This process was guided by the principle that categories should closely match users’ specific conditions. After this step, we have nine categories of tags and assign posts to each category based on their tags. To further improve the quality of the used posts in CanAnswer, we process the posts with user interaction data, i.e., the number of likes, comments, shares, collections, which were aggregated into a total engagement score. Guided by previous findings (Dobrian et al., 2011; Aiello et al., 2017; Noguti, 2016) that higher engagement correlates with higher quality, we selected the top 30 posts in each of the nine categories based on these scores. Finally, we manually reviewed the posts to exclude any remaining subtle advertisements, resulting in 246 posts across 9 categories, with each category containing between 8 and 30 posts. This comprehensive processing approach could enable users to receive high-quality, contextually relevant content that aligns closely with their specific needs and concerns.
@@ 6.1.3. Task and Procedure
We randomly assigned participants to a controlled group under baseline condition and a experimental group using CanAnswer. Initially, participants from baseline and CanAnswer groups were introduced to the study’s background, completed a screening survey, and reviewed the consent form. Upon consenting, participants received an introduction to their respective tools. Before beginning the tasks, we followed the approach of previous studies (Peng et al., 2019) and presented a scenario helping participants relate to a realistic situation in which someone might seek information about colorectal cancer. Two versions of scenario backgrounds were prepared ( shown in Appendix B), tailored to align with the participant’s age (under or above 35), representing people with early symptoms of colorectal cancer and different colorectal cancer risk factors who may need to learn about colorectal cancer (Patel and Ahnen, 2018; Venugopal and Stoffel, 2019; Saraiva et al., 2023; Johnson et al., 2013). These scenarios included details such as dietary habits, previous medical history, family health history, and specific symptoms that could prompt an individual to learn about colorectal cancer. Each participant was assigned to the corresponding scenario according to their age. Based on a pilot study with two participants, we suggested 30 minutes for the participants to engage with the provided scenario and learn about colorectal cancer using their respective tools (Figure 4), while informing them that they could adjust the duration as needed. After the learning session, we asked participants to verbally recall what they have learned and complete a knowledge test about colorectal cancer. They then filled out the questionnaires that evaluated their used system and joined a semi-structured interview to make sense of the ratings. On average, each participant spent 1 hour in the study and received 60 RMB as compensation.
@@ 6.1.4. Measurement
Learning Outcome. We asked the participants to verbally recall what they had learned about colorectal cancer after the learning session. Two authors first refined the coding scheme using four sampled transcripts, then independently coded ten documents each from the remaining set, and finally reconciled any differences. For each task we counted the number of correctly recalled knowledge points. Besides, we consulted four medical students (E2, E3 , E4, E5 in Table 4) who also participated in the expert interviews to compile a colorectal cancer knowledge test referred to (Green and Kelly, 2004). This Test  [^18] Contents are provided in the supplementary materials. includes seven single-choice questions ( five points per question), twelve true-or-false questions ( five points per question), and one scenario-based question ( twenty points), with a total 115 of points. In addition, we evaluate participants’ satisfaction, perceived usefulness, and reliability of the information gained in each condition using three 7-point Likert scale items adapted from (Peng et al., 2024, 2021, 2019).
Learning Experience. We measure the participants’ engagement in the learning session with six 7-point Likert scale items (Cronbach’s $\alpha=0.67$; 1 / 7 - strongly disagree/agree) adapted from (Peng et al., 2024; O’Brien, 2016). The items covered concentration, a sense of ecstasy, doability, a sense of serenity, timeless feeling, and intrinsic motivation. To measure the participants’ task workload, we adapted five questions from (Hart and Staveland, 1988; Fan et al., 2024), which are mental demand, physical demand, confusion, frustration, and cognitive load.
Usability of CanAnswer. Following (Fan et al., 2024), we measure the usability of CanAnswer with the ten questions from the System Usability Scale (SUS) (Brooke, 2013). In the interview with participants in the CanAnswer group, we asked their opinions on its key features.
@@ 7. Analyses and Results
This section presents quantitative and qualitative results from the evaluation studies. For the participants’ performance in the number of recalled knowledge points and score on the colorectal cancer knowledge test as well as the rated items about the learning experiences and perceptions of the learning tools, we performed a Mann-Whitney U tests (Mann and Whitney, 1947) to compare differences between the two groups. Two authors transcribed the recorded audio from both the between-subjects study and expert interviews into text scripts and conducted a thematic analysis of the data. They initially familiarized themselves with the transcripts by independently reviewing the text and coding the textual data. They then engaged in multiple rounds of comparison and discussion to refine the codes and ultimately finalized the coding scheme for all interview data. The occurrences of each code in the between-subjects study were counted (shown in Table 5), and these qualitative findings were incorporated into the following presentation of the results for the aforementioned measurements.

### v2

@@ 3.2. Design of Potential Features
Our design is informed by Kuhlthau’s Information Search Process (ISP) model (case2016looking; kuhlthau1991inside), which has been widely applied in educational contexts to understand how individuals seek and process information when faced with knowledge gaps. It outlines six stages that users typically go through in learning tasks: initiation, selection, exploration, formulation, collection, and presentation (kuhlthau1991inside). Based on this model and prior literature in the education domain, we designed six potential features to support users in learning health knowledge using data from online communities. For example, in the “Selection” stage of the ISP model, users are expected to identify and choose a general topic for investigation. However, grounded in the help_seeking’s research, individuals often experience difficulty initiating the search process due to uncertainty or lack of clarity about the topic. To address this, and inspired by the chunking mechanism in learning (gobet2001chunking), we designed a feature—Switch example questions about other topics—which categorizes community data into distinct topics, and we expected it could enable users to explore different directions and more easily select a focus for further learning. Similarly, in alignment with the “Exploration” stage, users are expected to investigate the content of a general topic to deepen their personal understanding. However, they may encounter difficulties in articulating their information needs precisely (kuhlthau1991inside; help_seeking). Inspired by prior tools that leverage LLMs to provide recommended questions, we designed the feature “Provide suggested follow-up questions” to support users in formulating relevant queries more easily during this process. Table 2 shows the six potential features, and the screenshots of these features are attached in Appendix A.
@@ 3.4. Findings
Provide suggested follow-up questions (help_seeking; pre-consultation; DesignQuizzer; wu2024comviewer) 6.01 1.00
Switch example questions about other topics (help_seeking; gobet2001chunking) 5.96 1.01
Offer input suggestions when the user enters the keyword (help_seeking; wu2024comviewer) 5.86 1.18
Provide real-world cases for reference (help_seeking; lieb2024student) 5.75 1.30
Assume a scenario to ask for an explanation (Learning_by_Teaching; lieb2024student; pan2025tutorup) 4.84 1.53
Test learning feedback with single-choice question (StayFocused; lieb2024student; DesignQuizzer; ruan2019quizbot) 4.64 1.87
We use the reflexive thematic analysis method (Thematic_analysis) to analyze the open-ended responses from the survey study as well as the transcribed recordings of each participant’s semi-structured interviews and co-design sessions.
@@ 3.5. Design Goals
DG1: The CA should ground its response to the user query on professional documents (fan2024lessonplanner; huatuo; adoption_decision). Authority is the first priority for knowledge acquisition in the health domain. Many online information, while easily accessible, could be unauthoritative (C1). We should prepare professional documents that are validated by experts (e.g., certified doctors) as the base for the tool to provide responses to the user query.
DG2: The CA should assist users in exploring multiple knowledge aspects of the disease (czeresnia1999concept; wicks2008patients), such as the cause, symbols, diagnosis, and treatment. This design goal highlights the breadth of knowledge coverage, aiming to help users develop a holistic understanding of the disease. The rich information online can cover every knowledge aspect, but such information is usually unstructured and non-personalized (C1). Categorizing frequently asked questions into topics could support users in exploring the knowledge aspects of their interests, and this feature was also anticipated by our participants ($M=5.96$).
DG3: The CA should provide real-world cases of others’ experiences related to the user’s (post_diversity; rare_disease; patientinfluencer). Participants went to online communities like RED to gain health information from others’ experiences. They can gain emotional support at the same time, which could reduce their anxiety (C2). The perceived usefulness of this feature was also highly rated by our participants ($M=5.75$).
DG4: The CA should assist users in delving into the current knowledge aspect of the disease (czeresnia1999concept; wicks2008patients). That is, it should support users in deepening their understanding of a specific topic they are currently exploring. This goal emphasizes the depth of knowledge acquisition and sustained engagement within a focused knowledge thread. Participants tended to give up seeking information online if they lacked support when confusion arose (C2). Suggesting follow-up questions ($M=6.01$) or input suggestions ($M=5.86$) could help them proceed to the learning process with CanAnswer, which were also favored by the participants in the study (Table 2).
@@ 5.1.1. Leveraging Doctor-Led Community Data
To satisfy DG1 and provide reliable, expert-verified responses, we first incorporate professional documents. Insights from our formative study suggest that tools designed to support health-related learning must ground their responses in trustworthy knowledge sources (fan2024lessonplanner; huatuo; adoption_decision). One potential approach is to curate a base dataset of professional documents (e.g., clinical guidelines, academic literature) with the assistance of domain experts. While such manually curated datasets can provide high-level guidance—such as helping determine appropriate topic categories for topic switching (DG4)—they are often insufficient in scope and depth for directly supporting large language models (LLMs) in generating trustworthy and detailed answers.
@@ 5.2.1. The Base Dataset of Professional Documents
To start with, we closely collaborated with a professional medical team to compile a database of clinical question-answering (QA) pairs related to colorectal cancer. First, we distributed a questionnaire to 68 staff in a local governmental project for screening colorectal cancer to collect the related questions commonly asked by people. We identified 16 frequently asked questions from the collected data. Second, we analyzed 100 publicly available doctor-patient conversation records from the Dingxiang Doctor, using the “colorectal cancer” and “colonoscopy” as the search keywords. This analysis yielded an additional 14 common questions. Third, we prompt GPT-4 to simulate queries by inputting the prompt, “If you were a first-degree relative of a colorectal cancer patient, what would you want to know about colorectal cancer and colonoscopy? ” This simulation identified one additional question that was not covered in the previous two sources. After further discussion within the medical team, our final base professional documents consist of 24 questions. The medical team compiles answers to these questions by referring to fifteen professional guidebooks  [^10] Full list of the guidebooks are provided in the supplementary materials. (e.g., “Asia Pacific Consensus recommendations for colorectal cancer screening” (sung2008asia)). To ensure the reliability of the compiled QA pairs, we consulted a panel of experts, including one physician specializing in cancer screening, two chief nurses with extensive experience in colorectal cancer care and colonoscopy screening, and five researchers specializing in cancer screening and disease prevention and control. The resulting 24 clinical QA pairs serve as the base dataset of professional documents in CanAnswer.
@@ 5.2.5. Topic switch
We utilize the 24 QA pairs in the Base Dataset ( Section 5.2.1) to group the topics related to colorectal cancers. The Enriched Dataset ( Section 5.2.2) was excluded from this step because it contains content, while relevant, that is not specifically about colorectal cancer. Given that the topics here aim to provide a broad, guideline-level overview specifically for learning about colorectal cancer, we opted to use only the Base Dataset. The initial categorization criteria were generated using LLOOM (lam2024concept), powered by LLMs, which creates 16 topics along with the criteria for classification. These topics were then reviewed and named by an expert. The finalized list of topics includes: Medical Definitions, Medication Use, Medical Decision-Making, Colon Cancer Treatment, Symptoms and Signs, Misconceptions About Colonoscopy, Family History Risks, Cancer Prevention, Dietary Focus, Risk Factor Identification, Pain Management, Follow-Up Care, Infection Prevention, Dietary Preparation, Colonoscopy Information, and Colon Cancer Screening Guidelines. When users interact with the topic switch dropdown menu, our system classifies their most recent query using GPT-4o, guided by refined standards generated by LLOOM. This mechanism accurately matches the user’s latest question to the most relevant topic in 16 topics. Once the user selects a new topic, CanAnswer prompts GPT-4o to generate follow-up questions by utilizing the chosen topic and the user’s chat history.
@@ 5.2.6. Processing Peer-led Online Community Data for CanAnswer
Our system recommends community posts from RED to provide users with relevant real-world cases related to colorectal cancer. We leveraged an open-source GitHub project  [^17] https://github.com/NanmiCoder/MediaCrawler to collect an initial dataset of 1,961 posts using keywords such as “intestinal polyps”, “hematochezia”, “intestinal cancer”, “chemotherapy regimen for colorectal cancer”, “immunotherapy for colorectal cancer”, “surgical resection for colorectal cancer”, “intestinal obstruction”, “colonoscopy”, “enteritis”, “flatulence”, and “diarrhea”. Then, we use tags associated with each post to classify these posts, as these tags were added by the post creators and could reflect the intentions and targeted audience of the posts. Specifically, these tags were input into GPT-4o, which generated a preliminary set of categories. The categories were iteratively refined through six rounds of feedback and manual verifications by two researchers, prioritizing the specificity of conditions and disease types to improve relevance. An expert (doctor with 30 years of clinical experience) then reviewed the final classification to validate its appropriateness. This process was guided by the principle that categories should closely match users’ specific conditions. After this step, we have nine categories of tags and assign posts to each category based on their tags. To further improve the quality of the used posts in CanAnswer, we process the posts with user interaction data, i.e., the number of likes, comments, shares, collections, which were aggregated into a total engagement score. Guided by previous findings (dobrian2011understanding; aiello2017beautiful; noguti2016post) that higher engagement correlates with higher quality, we selected the top 30 posts in each of the nine categories based on these scores. Finally, we manually reviewed the posts to exclude any remaining subtle advertisements, resulting in 246 posts across 9 categories, with each category containing between 8 and 30 posts. This comprehensive processing approach could enable users to receive high-quality, contextually relevant content that aligns closely with their specific needs and concerns.
@@ 6.1.3. Task and Procedure
We randomly assigned participants to a controlled group under baseline condition and a experimental group using CanAnswer. Initially, participants from baseline and CanAnswer groups were introduced to the study’s background, completed a screening survey, and reviewed the consent form. Upon consenting, participants received an introduction to their respective tools. Before beginning the tasks, we followed the approach of previous studies (peng2019design) and presented a scenario helping participants relate to a realistic situation in which someone might seek information about colorectal cancer. Two versions of scenario backgrounds were prepared ( shown in Appendix B), tailored to align with the participant’s age (under or above 35), representing people with early symptoms of colorectal cancer and different colorectal cancer risk factors who may need to learn about colorectal cancer (patel2018colorectal; venugopal2019colorectal; saraiva2023early; johnson2013meta). These scenarios included details such as dietary habits, previous medical history, family health history, and specific symptoms that could prompt an individual to learn about colorectal cancer. Each participant was assigned to the corresponding scenario according to their age. Based on a pilot study with two participants, we suggested 30 minutes for the participants to engage with the provided scenario and learn about colorectal cancer using their respective tools (Figure 4), while informing them that they could adjust the duration as needed. After the learning session, we asked participants to verbally recall what they have learned and complete a knowledge test about colorectal cancer. They then filled out the questionnaires that evaluated their used system and joined a semi-structured interview to make sense of the ratings. On average, each participant spent 1 hour in the study and received 60 RMB as compensation.
@@ 6.1.4. Measurement
Learning Outcome. We asked the participants to verbally recall what they had learned about colorectal cancer after the learning session. Two authors first refined the coding scheme using four sampled transcripts, then independently coded ten documents each from the remaining set, and finally reconciled any differences. For each task we counted the number of correctly recalled knowledge points. Besides, we consulted four medical students (E2, E3 , E4, E5 in Table 4) who also participated in the expert interviews to compile a colorectal cancer knowledge test referred to (CRC_test). This Test  [^18] Contents are provided in the supplementary materials. includes seven single-choice questions ( five points per question), twelve true-or-false questions ( five points per question), and one scenario-based question ( twenty points), with a total 115 of points. In addition, we evaluate participants’ satisfaction, perceived usefulness, and reliability of the information gained in each condition using three 7-point Likert scale items adapted from (DesignQuizzer; emotional_informational; peng2019design).
Learning Experience. We measure the participants’ engagement in the learning session with six 7-point Likert scale items (Cronbach’s $\alpha=0.67$; 1 / 7 - strongly disagree/agree) adapted from (DesignQuizzer; user_engagement). The items covered concentration, a sense of ecstasy, doability, a sense of serenity, timeless feeling, and intrinsic motivation. To measure the participants’ task workload, we adapted five questions from (NASA_TLX; fan2024lessonplanner), which are mental demand, physical demand, confusion, frustration, and cognitive load.
Usability of CanAnswer. Following (fan2024lessonplanner), we measure the usability of CanAnswer with the ten questions from the System Usability Scale (SUS) (SUS). In the interview with participants in the CanAnswer group, we asked their opinions on its key features.
@@ 7. Analyses and Results
This section presents quantitative and qualitative results from the evaluation studies. For the participants’ performance in the number of recalled knowledge points and score on the colorectal cancer knowledge test as well as the rated items about the learning experiences and perceptions of the learning tools, we performed a Mann-Whitney U tests (mann1947test) to compare differences between the two groups. Two authors transcribed the recorded audio from both the between-subjects study and expert interviews into text scripts and conducted a thematic analysis of the data. They initially familiarized themselves with the transcripts by independently reviewing the text and coding the textual data. They then engaged in multiple rounds of comparison and discussion to refine the codes and ultimately finalized the coding scheme for all interview data. The occurrences of each code in the between-subjects study were counted (shown in Table 5), and these qualitative findings were incorporated into the following presentation of the results for the aforementioned measurements.